In an 88 s interval, 588 hailstones strike a
glass window of area 0.624 m2
at an angle 20◦
to the window surface. Each hailstone has a
mass of 6 g and speed of 12.1 m/s.
If the collisions are elastic, find the average
force on the window.
Answer in units of N
Don't round answer.
1
Expert's answer
2020-02-05T12:15:41-0500

The impulse acting on a window


"\Delta p=2Nmv\sin 20^{\circ}"

The average force


"F=\frac{\Delta p}{\Delta t}=\frac{2Nmv\sin 20^{\circ}}{\Delta t}"


"=\frac{2\times 588\times 0.006 \times 12.1\times \sin 20^{\circ}}{88}=0.027\:\rm N=27\: mN."